# Designing Simple Mechanisms

### v1

@@ 4 Conclusion
Mechanisms can be complicated in many ways; challenging players to think contingently about other players’ moves, to plan for their own future moves, and to reason about other players’ beliefs. Thus, there are multiple ways to define what makes a ‘simple mechanism’. The right criterion depends on context; participants in designed mechanisms include schoolchildren, doctors, lumberjacks, fishermen, and telecommunications firms advised by game theorists (Athey and Levin, 2001; Marszalec et al., 2020; Bulow et al., 2009).
@@
Why have formal simplicity criteria at all? After all, we know it when we see it. If we use our intuition to judge formal criteria, why not use intuition instead of formal criteria? I see two considerations that weigh against the intuitive approach. First, simplicity is not the only goal—there are often others, such as efficiency, fairness, or revenue. Unless we put simplicity on an equal mathematical footing, we cannot study these trade-offs systematically. Without formal simplicity criteria, we may fixate on other desiderata that are well-formalized. Second, our intuitive judgements about simplicity are distorted by economic training. Experts in mechanism design suffer from the curse of knowledge; it is hard to adopt the perspective of people who do not know what we know (Camerer et al., 1989). In order to produce scientific knowledge about simplicity, we must construct precise theories that can be tested with data.
@@
There is a pressing need for experiments, both to test current theories about simplicity and to discover empirical regularities that future theories might explain. In testing theories, bear in mind that some tasks might be easy if done in isolation, but difficult when they are part of a broader strategic context. [^18] Moreover, in real-world mechanisms, participants must figure out their own preferences; each must devote “efforts and attention to an appraisal of the value the article would have in his own hands” (Vickrey, 1961), a complication that is often removed in laboratory experiments. Kagel and Levin (2016) survey experiments on auctions and Hakimov and Kübler (2021) survey experiments on matching mechanisms.
@@
Much remains to be done. The criteria we have studied so far are not exhaustive; there are other dimensions on which mechanisms can be simple or complex. There could be better ways to formalize the same dimensions, that are more tractable, more closely track human behavior, or have firmer cognitive foundations. Moreover, simplicity comes in degrees, and it would be useful to compare the relative simplicity of different mechanisms, or to find ways to say that one mechanism is ‘as simple as possible’ given other constraints. Nagel and Saitto (2023) have recently made progress in this direction. Finally, another important direction is to study how to describe and explain mechanisms, to help participants see for themselves that the mechanism is incentive-compatible (Gonczarowski et al., 2023).

### v3

@@ 5 Conclusion
Mechanisms can be complicated in many ways; challenging players to think contingently about other players’ moves, to plan for their own future moves, and to reason about other players’ beliefs. Thus, there are multiple ways to define what makes a mechanism “simple.” The right criterion depends on context; after all, participants in designed mechanisms include schoolchildren, doctors, lumberjacks, fishermen, and telecommunications firms advised by game theorists (Athey and Levin, 2001; Marszalec et al., 2020; Bulow et al., 2009).
@@
Why have formal simplicity criteria at all? After all, we know simplicity when we see it. If we use intuition to judge formal criteria, why not use intuition instead of formal criteria? I see two considerations that weigh against the intuitive approach. First, simplicity is not the only goal—there are often others, such as efficiency, fairness, or revenue. Unless we put simplicity on equal mathematical footing, we cannot study these trade-offs systematically. Without formal criteria for simplicity, we may fixate on other desiderata that are well-formalized.
Second, our intuitive judgements about simplicity are distorted by economic training. Experts in mechanism design suffer from the curse of knowledge; it is hard to adopt the perspective of people who do not know what game theorists know (Camerer et al., 1989). In order to produce scientific knowledge about simplicity, we must construct theories that can be tested with data.
@@
There is a pressing need for experiments, both to test current theories about simplicity and to discover empirical regularities that future theories might explain. In testing theories, bear in mind that some tasks might be easy if done in isolation, but difficult when they are part of a broader strategic context. Kagel and Levin (2016) survey experiments on auctions and Hakimov and Kübler (2021) survey experiments on matching mechanisms.
@@
Much remains to be done. The criteria for simplicity discussed here are not exhaustive; there are other dimensions on which mechanisms can be simple or complex. There could be better ways to formalize the same dimensions, that are more tractable, that track human behavior more closely, or that have firmer cognitive foundations. Moreover, simplicity comes in degrees, and it would be useful to compare the relative simplicity of different mechanisms, or to find ways to say that one mechanism is “as simple as possible” given other constraints. Nagel and Saitto (2023) recently made progress in this direction.
Finally, another important direction for the simplicity literature is to study how to describe and explain mechanisms, to help participants see for themselves that the mechanism is incentive-compatible (Gonczarowski et al., 2023). Real-world mechanisms, such as spectrum auctions and the National Resident Matching Program, often include detailed advice for participants. The Israeli Psychology Master’s Match even provided participants with a general-audience lecture demonstrating that its mechanism was strategy-proof, but some participants were clearly not convinced (Hassidim et al., 2021). Few laboratory experiments have studied such advice, possibly due to concerns about experimenter demand effects (De Quidt et al., 2019). The experiment conducted by Masuda et al. (2022) is a rare exception.